supose an investment of 4,500 doubles in value every ten years how much is the investment worth after 50 years?
Answer by ewatrrr(24785)


Hi
A= p(1+r/n) ^nt
A/p = 2 = (1+r) ^10
1.0718 = (1+r)
.0718 = r
A = $4500(1.0718)^50
A = $144,178.38